# EgoLifter: Open-world 3D Segmentation for Egocentric Perception

### v1

@@ Abstract
In this paper we present EgoLifter, a novel system that can automatically segment scenes captured from egocentric sensors into a complete decomposition of individual 3D objects. The system is specifically designed for egocentric data where scenes contain hundreds of objects captured from natural (non-scanning) motion. EgoLifter adopts 3D Gaussians as the underlying representation of 3D scenes and objects and uses segmentation masks from the Segment Anything Model (SAM) as weak supervision to learn flexible and promptable definitions of object instances free of any specific object taxonomy. To handle the challenge of dynamic objects in ego-centric videos, we design a transient prediction module that learns to filter out dynamic objects in the 3D reconstruction. The result is a fully automatic pipeline that is able to reconstruct 3D object instances as collections of 3D Gaussians that collectively compose the entire scene. We created a new benchmark on the Aria Digital Twin dataset that quantitatively demonstrates its state-of-the-art performance in open-world 3D segmentation from natural egocentric input. We run EgoLifter on various egocentric activity datasets which shows the promise of the method for 3D egocentric perception at scale. Please visit https://egolifter.github.io/ for video demonstrations and results.
@@ 1 Introduction
The rise of personal wearable devices has led to the increased importance of egocentric machine perception algorithms capable of understanding the physical 3D world around the user. Egocentric videos directly reflect the way humans see the world and contain important information about the physical surroundings and how the human user interacts with them. The specific characteristics of egocentric motion, however, present challenges for 3D computer vision and machine perception algorithms. Unlike datasets captured with deliberate "scanning" motions, egocentric videos are not guaranteed to provide complete coverage of the scene. This makes reconstruction challenging due to limited or missing multi-view observations.
The specific content found in egocentric videos also presents challenges to conventional reconstruction and perception algorithms. An average adult interacts with hundreds of different objects many thousands of times per day [4]. Egocentric videos capturing this frequent human-object interaction thus contain a huge amount of dynamic motion with challenging occlusions. A system capable of providing useful scene understanding from egocentric data must therefore be able to recognize hundreds of different objects while being robust to sparse and rapid dynamics.
@@ 2.2 Open-world 3D Segmentation
Recent research on open-world 3D segmentation [22, 14, 36, 50, 21, 42, 48, 17, 13, 43, 8, 29] has focused on lifting outputs from 2D open-world models - large, powerful models that are trained on Internet-scale datasets and can generalize to a wide range of concepts [38, 33, 20, 40]. These approaches transfer the ability of powerful 2D models to 3D, require no training on 3D models, and alleviate the need for large-scale 3D datasets that are expensive to collect. Early work [36, 14, 17] lifts dense 2D feature maps to 3D representations by multi-view feature fusion, where each position in 3D is associated with a feature vector. This allows queries in fine granularity over 3D space, but it also incurs high memory usage. Other work [11, 46, 24] builds object-decomposed 3D maps using 2D open-world detection or segmentation models [23, 20], where each 3D object is reconstructed separately and has a single feature vector. This approach provides structured 3D scene understanding in the form of object maps or scene graphs but the scene decomposition is predefined and the granularity does not vary according to the query at inference time. Recently, another work [1] lifts 2D instance segmentation to 3D by contrastive learning. It augments NeRF [31] with an extra feature map output and optimizes it such that pixels belonging to the same 2D segmentation mask are pulled closer and otherwise pushed apart. In this way, multi-view association of 2D segmentation is solved in an implicit manner and the resulting feature map allows instance segmentation by either user queries or clustering algorithms.
@@
We briefly review several recent and unpublished pre-prints that further explore topics in this direction using techniques similar to ours. Concurrently, OmniSeg3D [59] and GARField [18] follow the idea of [1], and focus on learning 3D hierarchical segmentation. They both take advantage of the multi-scale outputs from SAM [20] and incorporate the scales into the lifted features. GaussianGrouping [57] also approaches the open-world 3D segmentation problem but they rely on a 2D video object tracker for multi-view association instead of directly using 2D segmentation via contrastive learning. Similar to our improvement on 3DGS, FMGS [61] and LangSplat [37] also augment 3DGS with feature rendering. They learn to embed the dense features from foundation models [38, 34] into 3DGS such that the 3D scenes can be segmented by language queries. While the concurrent work collectively also achieves 3D reconstruction with the open-world segmentation ability, EgoLifter is the first to explicitly handle the dynamic objects that are commonly present in the real-world and especially in egocentric videos. We demonstrate this is a challenge in real-world scenarios and show improvements on it brought by EgoLifter.
@@ 3.3 Transient Prediction for Egocentric 3D Reconstruction
where the pixels with lower transient probability will contribute more to the reconstruction loss. As most of the objects in egocentric videos remain static, we also apply an $L$-1 regularization loss on the predicted $\mathbf{P}_{j}$ as $\mathcal{L}_{\text{reg}}(\mathbf{P}_{j})=\sum_{p\in\mathbf{P}_{j}}|p|.$ This regularization also helps avoid the trivial solution where $\mathbf{P}_{j}$ equals zero and all pixels are considered transient. The transient mask $\mathbf{P}_{j}$ is also used to guide contrastive learning for lifting instance segmentation, where the pixel set $\mathcal{U}$ is only sampled on pixels with the probability of being transient less than a threshold $\delta$. As shown in Fig. 2 and Fig. 3, this transient filtering also helps learn cleaner instance features and thus better segmentation results.
@@ 4.1 Benchmark Setup on ADT
4.2 Quantitative Results on ADT
4.3 Qualitative Results on Diverse Egocentric Datasets
In Fig. 3, we visualize the qualitative results on several egocentric datasets [35, 26, 10]. Please refer to project webpage [^1] https://egolifter.github.io/ for the videos rendered by EgoLifter with comparison to baselines. As shown in Fig. 3, without transient prediction, EgoLifter-Static creates 3D Gaussians to overfit the dynamic observations in some training views. However, since dynamic objects are not geometrically consistent and may be at a different location in other views, these 3D Gaussians become floaters that explain dynamics in a ghostly way, harming both the rendering and segmentation quality. In contrast, EgoLifter correctly identifies the dynamic objects in each image using transient prediction and filters them out in the reconstruction. The resulting cleaner reconstruction leads to better results in novel view synthesis and segmentation, as we have already seen quantitatively in Sec. 4.2. We also compare the qualitative results with Gaussian Grouping [57] in Fig. 4, from which we can see that Gaussian Grouping not only struggles with floaters associated with transient objects but also has a less clean feature map even on the static region. We hypothesize this is because our contrastive loss helps learn more cohesive identity features than the classification loss used in [57]. This also explains why EgoLifter-Static significantly outperforms Gaussian Grouping in segmentation metrics as shown in Tab. 1 and 2.
@@ 4.3 Qualitative Results on Diverse Egocentric Datasets
4.5 Limitations
We observe the transient prediction module may mix the regions that are hard to reconstruct with transient objects. As shown in rows (4) and (5) of Fig. 3, the transient prediction module predicts a high probability for pixels on the windows, which have over-exposed pixels that are hard to be reconstructed from LDR images. In this case, EgoLifter learns to filter them out to improve reconstruction on that region. Besides, the performance of EgoLifter may also be dependent on the underlying 2D segmentation model. EgoLifter is not able to segment an object if the 2D model consistently fails on it.
5 Conclusion
@@
Please refer to the video submitted together with this PDF. The supplementary video contains:
@@
A2.3 ADT Dataset Benchmark
A2.3.1 Sequence selection
@@
A2.3.2 Subset Splitting
@@
For sequences that only have a recording from one pair of Aria glasses, the first $4/5$ of the video is considered as seen views and the rest are considered as novel ones. For sequences that have videos from two pairs, the video from one pair is considered as seen views and the other is considered as novel views. During training, every 1 out of 5 consecutive frames in the seen views are used for validation the rest 4 are used for training. The entire novel subset is hidden from training and solely used for evaluation. For evaluation on 2D instance segmentation, we uniformly sampled at most 200 frames from each subset for fast inference. The objects in each video sequence are also split into dynamic and static subsets, according to whether their GT object positions have changed by over $2$cm over the duration of each recording. Humans are always considered dynamic objects.
A2.4 Results on ADT Seen Subset
@@
For completeness, we also report the 2D instance segmentation and photometric results on the seen subset of ADT in Tab. A.1. Note that the frames used for evaluation in the seen subset are closer to those for training, and therefore these results mostly reflect how well the models overfit to the training viewpoints in each scene, rather than generalize to novel views. As we can see from Tab. A.1, EgoLifter outperforms the baselines in segmenting static objects using both in-view and cross-view queries. When both static and dynamic objects are considered (the “All” column), EgoLifter still achieves the best results in cross-view, which is a harder setting for open-world segmentation. EgoLifter also has the second place in the in-view setting.

### v2

@@ Abstract
In this paper we present EgoLifter, a novel system that can automatically segment scenes captured from egocentric sensors into a complete decomposition of individual 3D objects. The system is specifically designed for egocentric data where scenes contain hundreds of objects captured from natural (non-scanning) motion. EgoLifter adopts 3D Gaussians as the underlying representation of 3D scenes and objects and uses segmentation masks from the Segment Anything Model (SAM) as weak supervision to learn flexible and promptable definitions of object instances free of any specific object taxonomy. To handle the challenge of dynamic objects in ego-centric videos, we design a transient prediction module that learns to filter out dynamic objects in the 3D reconstruction. The result is a fully automatic pipeline that is able to reconstruct 3D object instances as collections of 3D Gaussians that collectively compose the entire scene. We created a new benchmark on the Aria Digital Twin dataset that quantitatively demonstrates its state-of-the-art performance in open-world 3D segmentation from natural egocentric input. We run EgoLifter on various egocentric activity datasets which shows the promise of the method for 3D egocentric perception at scale. Please visit project page at https://egolifter.github.io/.
@@ 1 Introduction
The rise of personal wearable devices has led to the increased importance of egocentric machine perception algorithms capable of understanding the physical 3D world around the user. Egocentric videos directly reflect the way humans see the world and contain important information about the physical surroundings and how the human user interacts with them. The specific characteristics of egocentric motion, however, present challenges for 3D computer vision and machine perception algorithms. Unlike datasets captured with deliberate "scanning" motions, egocentric videos are not guaranteed to provide complete coverage of the scene. This makes reconstruction challenging due to limited or missing multi-view observations. The specific content found in egocentric videos also presents challenges to conventional reconstruction and perception algorithms. An average adult interacts with hundreds of different objects many thousands of times per day [5]. Egocentric videos capturing this frequent human-object interaction thus contain a huge amount of dynamic motion with challenging occlusions. A system capable of providing useful scene understanding from egocentric data must therefore be able to recognize hundreds of different objects while being robust to sparse and rapid dynamics.
@@ 2.2 Open-world 3D Segmentation
Recent research on open-world 3D segmentation [24, 16, 38, 54, 23, 44, 52, 19, 15, 45, 9, 31, 13] has focused on lifting outputs from 2D open-world models - large, powerful models that are trained on Internet-scale datasets and can generalize to a wide range of concepts [40, 35, 22, 42]. These approaches transfer the ability of powerful 2D models to 3D, require no training on 3D models, and alleviate the need for large-scale 3D datasets that are expensive to collect. Early work [38, 16, 19] lifts dense 2D feature maps to 3D representations by multi-view feature fusion, where each position in 3D is associated with a feature vector. This allows queries in fine granularity over 3D space, but it also incurs high memory usage. Other work [12, 49, 26] builds object-decomposed 3D maps using 2D open-world detection or segmentation models [25, 22], where each 3D object is reconstructed separately and has a single feature vector. This approach provides structured 3D scene understanding in the form of object maps or scene graphs but the scene decomposition is predefined and the granularity does not vary according to the query at inference time. Recently, another work [2] lifts 2D instance segmentation to 3D by contrastive learning. It augments NeRF [33] with an extra feature map output and optimizes it such that pixels belonging to the same 2D segmentation mask are pulled closer and otherwise pushed apart. In this way, multi-view association of 2D segmentation is solved in an implicit manner and the resulting feature map allows instance segmentation by either user queries or clustering algorithms.
@@
We briefly review several recent and unpublished pre-prints that further explore topics in this direction using techniques similar to ours. Concurrently, OmniSeg3D [63] and GARField [20] follow the idea of [2], and focus on learning 3D hierarchical segmentation. They both take advantage of the multi-scale outputs from SAM [22] and incorporate the scales into the lifted features. GaussianGrouping [61] also approaches the open-world 3D segmentation problem but they rely on a 2D video object tracker for multi-view association instead of directly using 2D segmentation via contrastive learning. Similar to our improvement on 3DGS, FMGS [66], LangSplat [39] and Feature3DGS [64] also augment 3DGS with feature rendering. They learn to embed the dense features from foundation models [40, 36] into 3DGS such that the 3D scenes can be segmented by language queries. While the concurrent work collectively also achieves 3D reconstruction with the open-world segmentation ability, EgoLifter is the first to explicitly handle the dynamic objects that are commonly present in the real-world and especially in egocentric videos. We demonstrate this is a challenge in real-world scenarios and show improvements on it brought by EgoLifter.
@@ 3.3 Transient Prediction for Egocentric 3D Reconstruction
where the pixels with lower transient probability will contribute more to the reconstruction loss. As most of the objects in egocentric videos remain static, we also apply an $L$-1 regularization loss on the predicted $\mathbf{P}_{j}$ as $\mathcal{L}_{\text{reg}}(\mathbf{P}_{j})=\sum_{p\in\mathbf{P}_{j}}|p|.$ This regularization also helps avoid the trivial solution where $\mathbf{P}_{j}$ equals one and all pixels are considered transient. The transient mask $\mathbf{P}_{j}$ is also used to guide contrastive learning for lifting instance segmentation, where the pixel set $\mathcal{U}$ is only sampled on pixels with the probability of being transient less than a threshold $\delta$. As shown in Fig. 2 and Fig. 3, this transient filtering also helps learn cleaner instance features and thus better segmentation results.
@@ 4 Experiments
We also compare EgoLifter with NeRF-based methods [50, 63] on the ADT dataset and evaluate EgoLifter on non-egocentric datasets [47, 1]. Please see Appendix A3 for the results.
@@ 4.1 Benchmark Setup on ADT
4.2 Quantitative Results on ADT
4.3 Qualitative Results on Diverse Egocentric Datasets
In Fig. 3, we visualize the qualitative results on several egocentric datasets [37, 28, 11]. Please refer to project webpage for the videos rendered by EgoLifter with comparison to baselines. As shown in Fig. 3, without transient prediction, EgoLifter-Static creates 3D Gaussians to overfit the dynamic observations in some training views. However, since dynamic objects are not geometrically consistent and may be at a different location in other views, these 3D Gaussians become floaters that explain dynamics in a ghostly way, harming both the rendering and segmentation quality. In contrast, EgoLifter correctly identifies the dynamic objects in each image using transient prediction and filters them out in the reconstruction. The resulting cleaner reconstruction leads to better results in novel view synthesis and segmentation, as we have already seen quantitatively in Sec. 4.2. We also compare the qualitative results with Gaussian Grouping [61] in Fig. 4, from which we can see that Gaussian Grouping not only struggles with floaters associated with transient objects but also has a less clean feature map even on the static region. We hypothesize this is because our contrastive loss helps learn more cohesive identity features than the classification loss used in [61]. This also explains why EgoLifter-Static significantly outperforms Gaussian Grouping in segmentation metrics as shown in Tab. 1 and 2.
@@ 4.3 Qualitative Results on Diverse Egocentric Datasets
5 Conclusion and Limitation
@@
Limitations: We observe the transient prediction module may mix the regions that are hard to reconstruct with transient objects. As shown in rows (4) and (5) of Fig. 3, the transient prediction module predicts a high probability for pixels on the windows, which have over-exposed pixels that are hard to be reconstructed from LDR images. In this case, EgoLifter learns to filter them out to improve reconstruction on that region. Besides, the performance of EgoLifter may also be dependent on the underlying 2D segmentation model. EgoLifter is not able to segment an object if the 2D model consistently fails on it.
@@
Acknowledgments
The authors thank the Project Aria team for providing open-source support and Nickolas Charron for helping with the ADT dataset evaluation. The authors also thank Fangzhou Hong, Kevin QH Lin, Hyo Jin Kim, Lingni Ma, Lambert Mathias, Pierre Moulon, Richard Newcombe, Tianwei Shen, Nan Yang, and Wang Zhao for discussions and useful feedback over the course of this project.
@@ Acknowledgments
Please refer to videos on the project page  [^1] https://egolifter.github.io/, which contains:
@@
A2.3 Timing
Using one NVIDIA A100 (40GB), training EgoLifter on one ADT sequence takes around 130 minutes (training vanilla 3DGS takes around 100 minutes). For a trained EgoLifter model, rendering both the RGB image and the instance feature map of $1408\times 1408$ resolution runs at around 103 fps. If only RGB images are rendered, the speed goes to 158 fps. Note that we use a different implementation than the original 3DGS, where we made several changes like not caching images on GPU to enable training on large datasets, e.g. AEA and Ego-Exo4D.
A2.4 ADT Dataset Benchmark
A2.4.1 Sequence selection
@@
A2.5 Comparison with NeRF
Method Nerfacto EgoLifter PSNR (all) 17.22 20.28
Table A.2: Comparison to Nerfacto on the ADT dataset.
Method INGP-Big M-NeRF360 3DGS EgoLifter PSNR 25.59 27.69 27.21 27.26
Table A.3: Quantitative comparison on the MipNeRF 360 dataset.
Figure A.2: Qualitative results on ADT datasets. From left to right: GT image; Render by Nerfacto; Render by EgoLifter.
Figure A.3: Qualitative results on MipNeRF 360. From left to right: GT image; EgoLifter RGB render; EgoLifter feature map (PCA).
Method Office0 Office1 Office2 Office3 Office4 Room0 Room1 Room2 Mean MVSeg [34] 31.4 40.4 30.4 30.5 25.4 31.1 40.7 29.2 32.4 SA3D [3] 84.4 77.0 88.9 84.4 82.6 77.6 79.8 89.2 83.0 OmniSeg3D [63] 83.9 85.3 89.0 87.2 78.3 83.0 79.4 88.9 84.4 EgoLifter (Ours) 82.9 78.4 85.1 84.1 80.0 77.0 85.4 84.3 82.1
Table A.4: Instance Segmentation results (mean IoU) on Replica dataset.
Figure A.4: Qualitative result on Replica datasets. From left to right: GT image; OmniSeg3D feature map; EgoLifter feature map
A2.5.1 Subset Splitting
@@
For sequences that only have a recording from one pair of Aria glasses, the first $4/5$ of the video is considered as seen views and the rest are considered as novel ones. For sequences that have videos from two pairs, the video from one pair is considered as seen views and the other is considered as novel views. During training, every 1 out of 5 consecutive frames in the seen views are used for validation the remaining 4 are used for training. The entire novel subset is hidden from training and solely used for evaluation. For evaluation on 2D instance segmentation, we uniformly sampled at most 200 frames from each subset for fast inference. The objects in each video sequence are also split into dynamic and static subsets, according to whether their GT object positions have changed by over $2$cm over the duration of each recording. Humans are always considered dynamic objects.
Appendix A3 Additional Results
A3.1 Results on ADT Seen Subset
@@
For completeness, we also report the 2D instance segmentation and photometric results on the seen subset of ADT in Tab. A.1. Note that the frames used for evaluation in the seen subset are closer to those for training, and therefore these results mostly reflect how well the models overfit the training viewpoints in each scene, rather than generalize to novel views. As we can see from Tab. A.1, EgoLifter outperforms the baselines in segmenting static objects using both in-view and cross-view queries. When both static and dynamic objects are considered (the “All” column), EgoLifter still achieves the best results in cross-view, which is a harder setting for open-world segmentation. EgoLifter also has the second place in the in-view setting.
A3.2 NeRF on ADT Dataset
In Tab. A.2 and Fig. A.2, we compare EgoLifter with the (default) nerfacto model in Nerfstudio [50] on the ADT dataset. As we can see from Fig. A.2, although Nerfacto uses per-image appearance embeddings to filter out transient phenomena in reconstruction, it still fails on challenging egocentric datasets like ADT and results in many floaters in the rendering. Quantitatively, EgoLifter also outperforms as shown in Fig. A.2.
A3.3 Non-egocentric public benchmarks
Non-egocentric benchmarks (Replica, ScanNet, MipNeRF 360) use careful hand-held scanning motions and lack dynamic phenomena. Therefore, they do not reflect the full capability of EgoLifter. We evaluate EgoLifter on the MipNeRF 360 dataset [1] in Tab. A.3 and Fig. A.3, where we use EgoLifter-static variant as the scenes are all static. Due to the lack of GT segmentation masks, we provide qualitative results on learned instance features in Fig. A.3. As shown in Tab. A.3 and Fig. A.3, EgoLifter has a similar PSNR as the original 3DGS and learns clean instance features that distinguish different instances.
We also test EgoLifter on Replica [47] and compare to OmniSeg3D [63], a recent feature lifting method based on NeRF representation and contrastive learning [2]. We evaluate the instance segmentation task using the multi-view mask propagation protocol [34, 3, 63], where the GT mask from one view is used for computing reference instance features and masks on other view are computed based on the feature distance from the reference ones. We follow the evaluation protocol in [63] and use Eq. (11) in [63] for computing the similarity scores. Similar to the experiments on MipNeRF 360, we used EgoLifter-static as there is no dynamic content in Replica scenes.
We report the quantitative results (in mIoU) in Tab. A.4 and a qualitative example in Fig. A.4. From Tab. A.4, we can see that EgoLifter has similar performance with the state-of-the-art NeRF-based segmentation methods [3, 63] on the non-egocentric Replica dataset. From the qualitative example in Fig. A.4, we can see that EgoLifter also results in clean and sharp feature boundaries on Replica as contemporary work OmniSeg3D [63], which distinguish different object instances and even the parts within each object.
Appendix A4 Additional Discussion on Limitations
Due to form factor and power constraints, egocentric videos are often captured with more challenges. Due to rapid head motion and lighting condition changes in the egocentric videos, the images contain significant motion blur that causes challenges in recovering sharp reconstructions from them. This explains in part the blurry results shown in some of the reconstruction results by EgoLifter. We leave how to improve the reconstruction quality from egocentric videos for future work.